A $53,000 loan is taken out on a boat with the terms 3% APR for 36 months. How much are the monthly payments on this loan? OA perpetuity will pay $700 per year, starting five years after the perpetuity is purchased. What is the present value (PV) ofWhat is the real interest rate given a nominal rate of 8.5% and an inflation rate of 4.5%? OA. 3.8% ○ B. 4.6% OC. OD. 5.4% 6.
Answer #1

Monthly Payment = Loan Amount/PVAF(r%, n periods)

= 53,000/PVAF(0.25%, 36 periods)

= 53,000/34.386465

= $1,541.30

i.e. C.

Value of perpetuity after 4 years = 700/7% = $10,000

Value of perpetuity today = 10,000/(1.07)4

= $7,629

i.e. D

(1+Nominal rate) = (1+Real Rate)(1+Inflation rate)

(1+8.5%) = (1+Real Interest rate)(1+4.5%)

Real rate = 3.8%

i.e. A.